OpenStudy (megannicole51):

what is the series equation of e^(-x^2) I have the expansion: 1-(x^2/1!)+(x^4/2!)-(x^6/3!)+... I know it is an alternating series so it will start with (-1)^n but I am making note cards for my exam so I need to make sure I have the equation right! @agent0smith

OpenStudy (anonymous):

\[e ^{x}=1+\frac{ x }{ 1! }+\frac{ x ^{2} }{2! }+\frac{ x ^{3} }{ 3! }+..........\] \[put x=-x ^{2}\] you are doing correctly.

OpenStudy (megannicole51):

what is the sum equation?